# Enhanced DeepONet for 1-D consolidation operator learning: an architectural investigation

### v1

@@ 8 Conclusion
This study demonstrates the performance of the DeepONet for learning solution operators in the 1-D consolidation problem, focusing on systematically investigating different DeepONet architectures. Our results reveal that feeding $C_{v}$ into the trunk net (model 3), rather than the more typical branch net (models 1 and 2), substantially improves prediction accuracy by 6 to 12 times lower test MSE. The improved accuracy is supported by the analytical form of the PDE, where $C_{v}$ modulates time-dependent behavior in the basis function. In addition, we further enhance the DeepONet architecture in Model 3 using Fourier Feature Embedding (Model 4). This further reduces the error by 1.5 times that of model 3, achieving the highest accuracy among all models. The architecture in Model 4 also effectively mitigates the high errors shown in highly varying, early-time excess PWP behaviors in the other model, due to its enhanced capacity to represent sharp spatio-temporal features from Fourier feature embedding. These findings provide guidance on architecture design for operator learning in similar PDE-driven systems governed by system coefficients.
@@
The trained DeepONet models achieve a significant computational speedup compared to traditional numerical solvers used to generate training data and benchmarks. Although the absolute time differences are on the order of milliseconds for the 1D consolidation case, the relative speedup–ranging from 1.5 to 100 times faster–highlights DeepONet’s potential as an efficient surrogate model. This computational advantage could be particularly valuable in large-scale or high-dimensional geotechnical simulations, where conventional solvers incur high costs due to time-stepping and meshing overhead.
@@
Our results also show that the trained DeepONet model performance diminishes when extrapolating to out-of-distribution values of $C_{v}$ and the function spaces for the initial excess PWP condition. This suggests that the deployment of a DeepONet-based surrogate model requires careful PDE parameter selection of a training dataset, which, in turn, requires relevant geotechnical expert knowledge.
@@
Beyond these technical contributions, this work introduces DeepONet to the geotechnical engineering community as an effective surrogate modeling framework. Its ability to generalize across varying PDE parameters and initial conditions without retraining makes it especially attractive for sensitivity analysis, real-time prediction, design optimization, and inverse problems, where traditional numerical solvers or even physics-informed neural networks may fall short due to computational or formulation constraints.
9 Acknowledgments
@@
This material is based on work supported by the National Science Foundation (NSF) under Grant No. CMMI 2145092. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the NSF.

### v2

@@ 8 Conclusion
This study demonstrates the performance of the DeepONet for learning solution operators in consolidation problems. Focusing first on the 1D case, we systematically investigate different DeepONet architectures. Our results reveal that feeding $C_{v}$ into the trunk net (Model 3), rather than the more typical branch net (models 1 and 2), substantially improves prediction accuracy by 6 to 12 times lower test MSE. The improved accuracy is supported by the analytical form of the PDE, where $C_{v}$ modulates time-dependent behavior in the basis function. In addition, we further enhance the DeepONet architecture in Model 3 using Fourier Feature Embedding (Model 4). This further reduces the error by 1.5 times that of Model 3, achieving the highest accuracy among all models. The architecture in Model 4 also effectively mitigates the high errors shown in highly varying, early-time excess PWP behaviors in the other model, due to its enhanced capacity to represent sharp spatio-temporal features from Fourier feature embedding. These findings offer guidance for designing DeepONet architectures for operator learning in PDE-driven systems governed by physical coefficients. Finally, we extend Model 4 to the 3D consolidation problem and demonstrate its use in an uncertainty quantification example.
@@
The trained DeepONet models achieve a significant computational speedup compared to traditional numerical solvers used to generate training data and benchmarks. In the 3D problem, DeepONet generates full 3D fields over time in less than a second, showing about three orders of magnitude faster computation speed than the high-fidelity finite difference-based solver. Although the absolute time differences are on the order of milliseconds for the 1D consolidation case, the relative speedup,ranging from 1.5 to 100 times faster, is still meaningful. Leveraging this efficiency, we show a conceptual demonstration of DeepONet’s potential to accelerate uncertainty quantification in 3D consolidation. This computational advantage could be particularly valuable in large-scale or high-dimensional geotechnical simulations, where conventional solvers incur high costs due to time-stepping and meshing overhead.
@@
Our results also show that the trained DeepONet model performance diminishes when extrapolating to out-of-distribution values of $C_{v}$ and the function spaces for the initial excess PWP condition. This suggests that the deployment of a DeepONet-based surrogate model requires careful PDE parameter selection of a training dataset, which, in turn, requires relevant geotechnical expert knowledge. If the problem can be described with a full PDE form, physics-informed loss terms can be employed to mitigate these output distribution errors (e.g., wang2021learning_pideeponet). Alternatively, active learning frameworks (e.g., guan2022adaptive_active; tharwat2023survey_active) could also be considered.
@@
Beyond the highlighted technical contributions, this work introduces DeepONet to the geotechnical engineering community as an effective surrogate modeling framework. Its ability to generalize across varying PDE parameters and initial conditions without retraining makes it especially attractive for real-time prediction, uncertainty quantification, design optimization, and inverse problems, where traditional numerical solvers or even physics-informed neural networks may fall short due to computational or formulation constraints.
In terms of engineering applications, we believe that the DeepONet model implemented in this study has the potential to be integrated with modern sensing technologies to enable pore-pressure monitoring and subsurface characterization, thereby supporting a real-time monitoring framework for consolidation-related problems (e.g., embankment consolidation, wick-drain performance). Although the example presented in the manuscript is based on synthetic data, it demonstrates DeepONet’s efficiency. Future studies that collect dense pore-pressure data and monitor the evolution of consolidation over time would be ideal for further advancing and validating the framework envisioned in this study.
Additionally, while this study uses a constant $C_{v}$ as trunk input, spatial heterogeneity could be addressed in future efforts by treating the soil property field $C_{v}(x,y,z)$ as an additional input function to the Branch network. Finally, a potential pathway for experimental integration could involve transfer learning, where a model pre-trained on large-scale synthetic data is fine-tuned using sparse, high-fidelity field measurements, providing a multi-fidelity DeepONet (xu2024multi).
Acknowledgments
@@
This material is based on work supported by the National Science Foundation (NSF) under Grant No. CMMI 2145092. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the NSF. This work is also supported by the InnoCORE program of the Ministry of Science and ICT (N10250154).